OpenStudy (anonymous):

if the point P(x,2/3) is on the unit circle and P is in the 2nd quadrant, what is x equal to?

OpenStudy (anonymous):

now you are in the second quadrant, so what is the sign of the first coordinate?

OpenStudy (anonymous):

negative

OpenStudy (anonymous):



OpenStudy (anonymous):

ok good because you are two the left

OpenStudy (anonymous):

since you are on the unit circle, like last time you know that \[x^2+\left(\frac{2}{3}\right)^2=1\]

OpenStudy (anonymous):

with a little practice you can go right to \[x=\sqrt{1-\left(\frac{2}{3}\right)^2}\]

OpenStudy (anonymous):

unlike the last time, you will have a radical in your answer let me know what you get

OpenStudy (anonymous):

so x would equal either √(5)/3, or -√(5)/3 ?

OpenStudy (anonymous):

exactly and of course you know which one to pick right?

OpenStudy (anonymous):

the negative, correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

great! thanks so much :) you really helped a lot with this

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

you did all the work anyway